CH/DOC/92/01

List of Names

Distribution List

Compiled: March 10, 1993

Postal Addresses of Institutions

Name  Address  STD code  Fax
Birmingham School of Physics,  021  414-6709

University of Birmingham, Edgbaston, Birmingham B15 2TT.

Surrey  Department of Physics,  0483  304212

University of Surrey, Guildford, Surrey GU2 5XH.

York  Physics Department,  0904  432-214

University of York, Heslington, York YO1 5DD.

Oxford  Department of Nuclear Physics,  0865  273-418

University of Oxford, Keble Road, Oxford OX1 3RH.

Canberra  Department of Nuclear Physics,  010+61-6 249-0748

Research School of Physical Sciences, Institute for Advanced Study, The Australian National University, GPO Box 4, Canberra, ACT, Australia. 2601

continued…

Name  Address  STD code  Fax
Daresbury SERC Daresbury Laboratory,  0925  60-3173

Warrington,

Cheshire WA4 4AD.

Manchester Department of Physics,  061  273-5867

Schuster Laboratory, The University, Manchester, M13 9PL.

Academic Staff

Name  Electronic Mail Address  STD Telephone

Dr W.N. Catford  W.Catford@ph.surrey.ac.uk 0483 509323

Dr N.M. Clarke  NMC@np.ph.bham.ac.uk  021 414-4680
Dr B.R. Fulton  BW@np.ph.bham.ac.uk  021 414-4683
Dr J.S. Lilley  Lilley@Daresbury.ac.uk  0925 60-3558
Dr W.D.M. Rae  Rae%ph.ox@ph.ox.ac.uk  0865 273-368
Dr G. Tungate  PSI@np.ph.bham.ac.uk  021 414-4685
Dr D.L. Watson  Watson@phga.york.ac.uk  0904 432-245

Research Staff

Name  Electronic Mail Address  STD Telephone
Mr R.A. Cunningham RAC@nnga.daresbury.ac.uk 0925 60-3431

Dr M. Freer  MF@np.ph.bham.ac.uk  021 414-4685

Dr G.J. Gyapong  G.Gyapong@ph.surrey.ac.uk 0483 509405

Dr S.J.A. Hall  SJAH@np.ph.bham.ac.uk  021 414-4685

Mr R.A. Hunt  Hunt%ph.ox@ph.ox.ac.uk  0865 273-333

Mr C.D. Jones  Jones@phga.york.ac.uk  0904 432-245

Dr A.C. Merchant  Merchant%ph.ox@ph.ox.ac.uk 0865 273-312

Mr V. Pucknell  Pucknell@daresbury.ac.uk 0925 60-3431

Dr R.P. Ward  RPW@np.ph.bham.ac.uk  021 414-4685

Dr Jie Zhang  Zhang%ph.ox@ph.ox.ac.uk 0865 273-312

Research Students

Name  Electronic Mail Address  STD Telephone
Mr S.P. Chappell  OEF@phga.york.ac.uk  0904 432-245
Mr N. Curtis  NC@np.ph.bham.ac.uk  021 414-4683
Mr S.P. Fox  Fox@phga.york.ac.uk  0904 432-245
Mr M.J. Leddy  ML@np.ph.bham.ac.uk  021 414-4683
Mr A.St.J. Murphy  AM@np.ph.bham.ac.uk  021 414-4683
Ms J.S. Pople  JP@np.ph.bham.ac.uk  021 414-4683
Mr P.M. Simmons  Simmons%s10.ph.ox@ph.ox 0865 273-333
Mr S.M. Singer  S.Singer@surrey.ac.uk  0483 509405